OpenStudy (idealist10):

Find a curve y=y(x) through (1, 2) such that the tangent to the curve at any point (xo, y(xo)) intersects the x axis at xI=xo/2.

OpenStudy (abb0t):

@Luigi0210 ?

OpenStudy (luigi0210):

Sorry, don't know @satellite73 @ganeshie8 @Directrix

OpenStudy (anonymous):

eq. of any line through ( x0,y(x0)) is y-y(x0)=m(x-x0), where m is the slope. because it passes through\[(\frac{ xo }{ 2 },0)\] \[0-y(x0)=m(\frac{ x0 }{ 2 }-x0)\] \[-y(x _{0})=-m \frac{ x _{0} }{ 2 }\] \[\frac{ 2y \left( x _{0} \right) }{ x _{0} }=m ~at ~(x _{0},y(x _{0}))\] \[m=\frac{ dy }{ dx }=\frac{ 2y }{ x }\] separating the variables and integrating \[\int\limits \frac{ dy }{ y }=2\int\limits \frac{ dx }{ x }+C\] \[ \ln y=2 \ln x+C\] \[\ln y=\ln x^2+C\] \[\ln y-\ln x^2=C\] \[\ln \frac{ y }{ x^2 }=C\] \[\frac{ y }{ x^2 }=e^C=c (say)\] \(because~ it~ passes ~through~ (1,2)\) \[\frac{ 2 }{ 1^2 }=c,c=2\] ?

OpenStudy (idealist10):

Thank you!

OpenStudy (anonymous):

yw